A motorcycle is traveling up one side of a hill and down the other side. The crest of the hill is a circular arc with a radius of 49.4 m. Determine the maximum speed that the cycle can have while moving over the crest without losing contact with the road.
Answer #1

V = SQRT [ AR ]

V = Velocity
A = Acceleration
R = Radius

Given


A = 9.80665 m/s^2
R = 49.4 m

V = SQRT [ (9.80665 m/s^2) * (49.4 m) ]
V = SQRT [ 484.12 m^2/s^2 ]
V = 22.00 m/s
Answer #2

Fn = Fg = mg = Fc = mv^2/r
g = v^2/r
v^2 = gr
v = sqrt(gr)
Answer #3

V = SQRT [ AR ]

V = Velocity
A = Acceleration
R = Radius

Given


A = 9.80665 m/s^2
R = 49.4 m

V = SQRT [ (9.80665 m/s^2) * (49.4 m) ]
V = SQRT [ 484.44 m^2/s^2 ]
V = 21.8 m/s